a file cabinet manufacturer estimates that 5% of its file cabinets will have drawers that close improperly. Assume a probably run of 40 cabinets is completed. a.what is the mean and standard deviation of this distribution? b. what is the probability that 3 or more are defective? c. what is the probability that exactly 3 cabinets are defective? d. what is the probability that there are more than 4 defective file cabinets?


1
Expert's answer
2021-07-12T16:16:18-0400

Let "X=" the number of defective file cabinets: "X\sim Bin(n, p)."

Given "n=40, p=0.05, q=1-p=1-0.05=0.95."

a.


"\mu=np=40(0.05)=2"

"\sigma^2=npq=40(0.05)(0.95)=1.9"

"\sigma=\sqrt{\sigma^2}=\sqrt{1.9}\approx1.3784"

b.


"P(X\geq 3)=1-P(X=0)-P(X=1)"

"-P(X=2)=1-\dbinom{40}{0}(0.05)^0(0.95)^{40-0}"

"-\dbinom{40}{1}(0.05)^1(0.95)^{40-1}-\dbinom{40}{2}(0.05)^2(0.95)^{40-2}"

"\approx0.32326423925"


c.


"P(X= 3)=\dbinom{40}{3}(0.05)^3(0.95)^{40-3}"

"\approx0.18511446375"

d.


"P(X>4)=1-P(X=0)-P(X=1)"

"-P(X=2)-P(X=3)-P(X=4)"

"=1-\dbinom{40}{0}(0.05)^0(0.95)^{40-0}"

"-\dbinom{40}{1}(0.05)^1(0.95)^{40-1}-\dbinom{40}{2}(0.05)^2(0.95)^{40-2}"

"-\dbinom{40}{3}(0.05)^3(0.95)^{40-3}-\dbinom{40}{4}(0.05)^4(0.95)^{40-4}"

"\approx0.04802826025"